The total height of a statue ans its pedestal is 305 feet. If the statue is 3 feet shorter than the height of the pedestal, find the height of the statue.
Answer by nyc_function(2741)

Let (x - 3) = height of statue

Let x = height of pedestal


Total length = 305



x + (x - 3) = 305


2x - 3 = 305


2x = 305 + 3


2x = 308


x = 308/2


x = 154


Height of statute = x - 3.


We know that x = 154.


Height of statute = 154 - 3


Height of statute = 151 feet.